How many grams of lead (II) oxide will be produced by the decomposition of 2.50 g of lead (II) carbonate? *
1
Expert's answer
2020-03-27T08:59:17-0400

The balanced equation for considered reaction can be written as follows

"PbCO_3 \xrightarrow{t \degree C} PbO + CO_2"


"\frac{m(PbCO_3)}{M(PbCO_3) \cdot n(PbCO_3)} = \frac{m(PbO)}{M(PbO) \cdot n(PbO)}",

where "m" - mass, g; "M" - molar mass, g/mol; "n" - number of molecules in the balanced equation.

"m(PbO) = \frac{m(PbCO_3) \cdot M(PbO) \cdot n(PbO)}{M(PbCO_3) \cdot n(PbCO_3)}"

"M(PbCO_3) = 1 \cdot 207.2 + 1 \cdot 12.011 + 3 \cdot 15.9994 \approx 267.21 \; g/mol"

"M(PbO) = 1 \cdot 207.2 1 \cdot 15.9994 \approx 223.21 \; g/mol"

"m(PbO) = \frac{2.50 \cdot 223.21 \cdot 1}{267.21 \cdot 1} \approx 2.09 \; g"


Answer: 2.09 g.